Fermi National Laboratory


Accelerator Update

Friday April 16
The day shift began with Tevatron (TeV) experts conducting studies until the repair of the abort kicker (AAK1) begins, with Recycler experts conducting studies, with the Antiproton Source (Pbar) stacking, and with MiniBooNE and the Meson MCenter experiment taking beam.

At 8:40 AM, Operations turned off beam to Switchyard (SY).

Operations turned off the TeV to begin work on AAK1.

Operations set up the Recycler to stash antiprotons from Pbar at 10:56 AM. The stashing ended at 1:08 PM. Pbar resumed stacking.

Operators established beam to the SY dump at 2:22 PM, and then established beam to MTest at 3:07 PM.

At 3:44 PM, a TeV expert reported that the kicker work had finished a while ago, but conditioning would continue.

Booster experts began a cogging study at 4:10 PM that required SY and MiniBooNE beam to be off and stacking halted. The study ended at 5:30 PM and beam resumed.

Operations put the TeV through a dry squeeze 6:31 PM, through a wet squeeze at 7:04 PM, and then TeV experts began studies.

At 9:38 PM, Operations put the TeV into shot setup.

At 11:53 PM, Operations reported that they had lost protons out of the first bunch injected into the TeV. They aborted the protons and checked chromaticity.

Saturday April 17
The midnight shift began with the TeV in shot setup, with Operations monitoring the Recycler stash, and with MiniBooNE taking beam.

At 1:07 AM, Operations established store 3411 with an initial luminosity of 44.55E30.

Operations established beam to the SY dump at 1:15 AM.

At 3:16 AM, Recycler began stashing more antiprotons from Pbar.

Pbar resumed stacking at 3:35 AM.

Operations reestablished beam to SY at 3:51 AM.

At 8:33 AM, the TeV suffered a sector F4 quench.

Operations prepared the Collider Detector at Fermilab (CDF) and D-zero (D0) for access at 9:44 AM.

At noon, Operations reported that the evidence suggested a TeV antiproton abort kicker (AAK1) prefire caused the quench. Operators called in experts who confirmed the prefire, followed by repairs.

CDF and D0 came out of access at 12:22 PM.

Operations established beam to MTest at 3:23 PM.

CDF went back in on access for an hour at 4:05 PM.

At 4:12 PM, Operations reported a problem with a Pbar Debuncher RF station (DRF1) alarming. Operators called an expert who believed the problem was with the power amplifier bias malfunctioning. This wasn't causing problems with stacking.

Operations put the TeV through a dry squeeze at 6:12 PM.

Operations put the TeV in shot setup at 7:14 PM

Operations reported at 8 PM that they just had the third beam abort in the Linac 400 MeV area due to a noisy output of the Booster chopper (B:CHOP). Operators called in an expert to investigate.

At 9:14 PM, Operations established store 3414 with an initial luminosity of 39.84E30.

Sunday April 18
The midnight shift began with Operations monitoring store 3414 and the Recycler stash, with Pbar stacking, and with MiniBooNE and the Meson experiments taking beam.

At 9:48 AM, a Main Injector (MI) expert began a barrier bucket study.

Operations terminated store 3414 at 7:13 PM.

Operators prepared CDF and D0 for an access at 7:30 PM. Both experiments were out of access by 9:21 PM.

Operations put the TeV into shot setup at 10:30 PM. Operators disabled beam to SY.

Cryo system experts began a heat-load test on the system at TeV sector F1 at 10:34 PM.

Monday April 19
The midnight shift began with the TeV shot setup on hold for a cryo system heat-load test at sector F1, with Operations monitoring a Recycler stash, and with MiniBooNE and MCenter taking beam.

At 12:03 AM, cryo system experts reported a problem with the sector F1 wet engine. They had the engine repaired by 1:40 AM.

Operations sent a pilot-shot of antiprotons from Pbar to the MI abort dump at 1:55 AM.

At 2:31 AM, Operations established store 3416 with an initial luminosity of 41.04E30. Pbar resumed stacking.

The TeV sector C17 separator started sparking at 3:40 AM. It continued to spark, luminosity began to drop, and beam started to fall out of the machine. CDF and D0 turned off the high voltage to their detectors.

Operations terminated store 3416 at 4:40 AM.

A TeV expert came in and put the C17 separator into conditioning mode at 5:53 AM.

Plans for Today
The plans for today are to allow TeV experts to investigate and resolve the separator problem and then establish a new store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/19/2004   email Fermilab